If you are rolling a fair die, what is the probability of rolling an even number and then, on a second roll, rolling the very same number that you rolled on the first roll.
(even# and same #)=

Answer by venugopalramana(3286)

If you are rolling a fair die, what is the probability of rolling an even number and then, on a second roll, rolling the very same number that you rolled on the first roll.
(even# and same #)=
TOTAL THERE ARE 6 FACES
THEY HAVE 1 TO 6 NUMBERS
EVEN NUMBERS ARE 3
SO PROBABILITY OF GETTING AN EVEN NUMBER ON I THROW = 3/6
TO GET SAME NUMBER IN II THROW PROBABILITY=1/6
HENCE COMBINATION PROBABILITY = (3/6)(1/6)=1/12